### v1

# The nature of ASASSN-24fw’s occultation: modelling the event as dimming by optically thick rings around a sub-stellar companion

Abstract
ASASSN-24fw, a main-sequence star monitored by the All-Sky Automated Survey for SuperNovae, faded rapidly in late 2024, with the event lasting until June 2025. The pre-dimming SED of ASASSN-24fw indicates that it is a F-type main sequence star with persistent infrared excess having a fractional luminosity of around 0.5%. We modeled this infrared excess using two-component blackbody models and find that the warmer component has a effective temperature of 1072K while the colder component has an effective temperature of 387K. The publicly available survey lightcurves show that ASASSN-24fw is a stable star in brightness and hence the dimming is due to external companion. While long period stellar dimming events have been increasingly discovered over the past decade, what distinguishes the 275 days long event ASASSN-24fw is its unique "flat-base" in the lightcurve, which has a duration of nearly 200 days. In order to discern the physical properties of the occulter and the host star, we have analyzed this event using available photometric and spectroscopic data taken during dimming. In particular, we have fitted two different lightcurve models and performed spectroscopic analysis. Our first modeling method quantifies the parameters of the lightcurve and reveals the multiple phases of ingress, thereby indicating variation in the density of the occulting material. Our second lightcurve model suggests that a gas-giant (perferably a brown dwarf) with a minimum mass of 3.42MJ and having massive (0.17 au) circum-planetary rings. The near-infrared spectra taken during dimming show elevated levels of infrared excess and also indicate that the central object is likely a brown dwarf with a spectral-type of M8. We notice that the H line is transient in nature and has relation with the composition of the occulter. We observed this event using LCOGT to measure the extinction during dimming. We find that the the occulting material may contain large bodies and gas. We also find a companion in the LCOGT images within 3" radius. Our analysis suggests that this is a cool main-sequence star at a separation of 3000 au and unlikely to be un-associated with ASASSN-24fw. Finally, we suggest that systems like ASASSN-24fw are very rare and unique and more follow-up observations are needed to constrain the stellar properties, the circum-stellar environment and the evolutionary mechanisms for such systems.
keywords:
binary – stellar dimming – circum-stellar disc – exoplanet1 Introduction
Nearly all stars are born surrounded by dense circum-stellar envelopes (Ribas et al., 2014; Manara et al., 2023). These envelopes process the natal proto-stellar cloud into planetary discs over a period of a few to 10s of million years (Williams & Cieza, 2011). These circum-stellar young protoplanetary discs are the birthplace of planetary systems including planetesimal belts (Wyatt et al., 2015; Lovell et al., 2021). Circum-stellar discs manifest in various forms (Hughes et al., 2018) and now we know that they are ubiquitous throughout the pre- to post-main sequence lifetimes of stars (Jewitt et al., 2009; Dennihy et al., 2020; Zhao et al., 2020; Raymond & Nesvorný, 2022; Myers et al., 2024; Swan et al., 2024; Farihi et al., 2025). The dynamical excitation induced erosive or destructive collisions of the planetesimals (e.g. a phase of Heavy-Bombardment period in Solar System (Gomes et al., 2005; Bottke & Norman, 2017)) is a common mechanism to generate detectable levels of dust manifesting as debris discs around stars (e.g. Marshall et al., 2021; Matrà et al., 2025).
Circum-stellar discs significantly change their structure, dynamics, and composition over time and, thus their architectures and brightness are correlated with the stellar age (Cao et al., 2023; Matrà et al., 2025), e.g., a pre-main sequence is surrounded by a protoplanetary disc of mainly gas and dust, the transition or intermediate disc phase has altered composition and geometry while the debris disc is further evolved like the star hosting it and contains mainly of collisional rocky bodies and dust generated by destructive processes (Hughes et al., 2018; Su et al., 2019; Moór et al., 2024). Thus, the detection of such debris discs points towards successful planet-formation (or even destruction) processes.
These circum-stellar discs can be detected through the presence of excess flux from infrared to millimeter wavelengths (Su et al., 2019). As the material in these discs re-emits the incident stellar radiation into longer wavelengths, the composition of the debris disc, structure and its evolutionary status can be inferred from the location of the peak of the excess flux (Wyatt et al., 2015; Hughes et al., 2018; Cao et al., 2023). This information in parallel with the nature of the optical variability can be used to understand the nature of the circum-stellar material e.g., either proto-planetary, transitional, or debris discs. For example, V488 Per has the most infrared excess of any circum-stellar disc and is therefore thought to be surrounded by an Extreme Debris Disc (Moór et al., 2024) which has a fractional luminosity 1%, HD 166191 is surrounded possibly by a transition disc (Kennedy et al., 2014) (although this is still controversial, Garcia & Hughes, 2019), and RZ Piscium contains a debris disc perturbed by a red-dwarf companion (Kennedy et al., 2020). However, there are examples of optical variability without infrared excess e.g. Boyajian’s star (KIC 8452642; Boyajian et al., 2016) had no transient infrared excess (Marengo et al., 2015) and therefore it was attributed to the disintegration of a family of exo-comets (Bodman & Quillen, 2016). Another such event around ASASSN-21qj has likewise been attribted to exocometary activity and exhibited a transient infrared excess (Marshall et al., 2023).
Sometimes, the variability in stellar brightness is also caused by systems containing circum-planetary, or circum-stellar discs. For example, the Mamajek’s object (Mamajek et al., 2012), RW Aur (Rodriguez et al., 2013; Rodriguez et al., 2016), V409 Tau and AA Tau (Rodriguez et al., 2015), "The Random Transitor" type objects HD 139139 (Rappaport et al., 2019b), and EPIC 204376071 (Rappaport et al., 2019a), Gaia21bcv (Hodapp et al., 2024), ASASSN-21js (Pramono et al., 2024). Another example is VVV-WIT-08/10 (Smith et al., 2021), discovered by the VISTA Variables in the Vía-Láctea (VVV; Minniti et al. (2010)) survey, where an eclipse lasting 200 days is attributed to an elliptical disc-like occulter, similar to the proposed ring system of J1407b (Kenworthy & Mamajek, 2015).
One of the important distinguishing points between variability due to circum-stellar material and circum-planetary material is the depth and duration of the dimming. While the variability due to the former is irregular and a combination of several short-period dips, the variability due to the latter could manifest as a dip both larger in depth and period. It may also show some extra features depending on the orientation of the system. As it is not exactly clear how such objects with massive rings form and if there is any correlation between the circum-stellar environment and the properties of such objects, we can build a clearer picture about the such systems only by continuing to identify and analyze such events.
This paper is another attempt to explain the recent stellar dimming event: ASASSN-24fw. This dimming event was reported by ASAS-SN’s transient detection page followed by an Astronomer’s Telegram (ATel No. 16833, JoHantgen et al., 2024). ASASSN-24fw has been the subject of study in two recently published articles by Forés-Toribio et al. (2025) and Zakamska et al. (2025). Our preliminary investigation found that ASASSN-24fw is located close to a star forming region in the Monoceros constellation (Hughes & Baines, 1985; Costado & Alfaro, 2018). However, later we found that both are not related to each other. The star has a persistent infrared excess with a fractional luminosity of 0.5%. We find that a two component blackbody model with effective temperatures of 1072K and 387K can explain the infrared excess in the SED. Since it had a stable optical and mid-infrared pre-dimming brightness, it can be assumed that the circum-stellar material is distributed around the star. In September 2024, it experienced a steep reduction in brightness which lasted until around the first week of October 2024. The star remained around the same low-brightness until the mid-May of 2025 and the egress occurred close to the first week on June 2025. Another Telegram by Nair & Denisenko (2024) (ATel No. 16919) noted similar dimming episodes of this star around the years 1937 and 1982, suggesting a possible periodicity of 16,000 days (or 43.8 years). The drastic drop in the brightness of ASASSN-24fw from g’ 13.20 mag to 16.70 mag (close to the limit of ASAS-SN sensitivity at 17.50 mag) rapidly within a fortnight followed by a nearly stable brightness for six months and mirrored egress is the unique feature of this event.
From the shape and analysis of the lightcurve, we suggest that the dimming actually started around March 2024 and the ingress happened in three stages. We suggest that the dimming was caused by a central body with massive circum-planetary rings ( 0.17 au in size). The central body is a gas-giant with a mass of at least 3.42MJ and the material density in the rings is inversely proportional to the radial distance from it. We think that the gas-giant is likely to be a brown-dwarf although we do not rule out a super-Jupiter interpretation. We present previously unpublished information tracing the evolution of this event after its discovery, seeking to characterize both the size and mass of the star and the occulting body, and identify its size, chemical composition, and mass of the constituent material around it. We also discuss the studies by Forés-Toribio et al. (2025) and (Zakamska et al., 2025) later in our paper from a broader perspective.
2 Data
2.1 All-Sky Automated Survey for SuperNovae (ASAS-SN)
The All-Sky Automated Survey for SuperNovae (ASAS-SN) (Shappee et al., 2014) is a long term survey project to monitor the sky down to a limiting magnitude of V 17 mag using a global network of telescopes. The focus of the survey is to find nearby supernovae (SNe) and other transient sources. Currently, this network consists of six fully robotic units on Mount Haleakala in Hawaii, Cerro Tololo in Chile, LCO sites in South Africa and USA, and another in China. Each unit consists of several robotic 14 cm aperture telescopes and is hosted by the corresponding station and each telescope is mounted with a Nikon telephoto lens and a 2k x 2k thinned CCD, giving a 4o.5 x 4o.5 field-of-view and a 7.8" pixel scale. These 24 telescopes together survey entire night sky every night and discoveries are announced within few hours of the data being taken. The data can be accessed through https://asas-sn.osu.edu/.
2.2 LCOGT
We monitored ASASSN-24fw with time-series photometry from February 22nd to April 5th 2025 (PID:DDT2025A-003, PI: Sarang Shah) using the Las Cumbres Observatory’s network of 1-m telescopes with the SINISTRO imaging instrument (Brown et al., 2013). The stellar brightness was recorded in four filter bands: SDSS g’, r’, i’, and PANSTARRS zs. Integration times were 12s in g’, 18s in r’, 26s in i’, and 94s in zs. This provided a SNR100 in each filter for an occulted magnitude of 16.5. The photometric measurements were taken from the pipeline-produced BANZAI catalog (McCully et al., 2018). If the catalog (or target) was missing from the observatory provided data, the stellar magnitude was determined using aperture photometry performed using the Python-based Photutils package (Bradley et al., 2024) using a 5" radius aperture and a background annulus between 10" and 15".
In the LCOGT images, we find a faint companion star which is also present in Gaia DR3 with > 19 mag and at a separation of 3". In Figure (1), we show a cropped 3"x3" LCOGT g-band image where the colorbar represents the pixel count. ASASSN-24fw and its faint companion are located in the center of the red-circle (and image). If bound to ASASSN-24fw, at a distance of 1kpc, this companion is located at a distance of 3000 au. To check how likely it is to have a chance alignment for a faint companion at a separation 3", we stacked our images to increase the SNR of the companion. We then performed the psf photometry of the companion by masking ASASN24fw. The (g’-r’), (r’-i’), and (i’-z’) colors that we get are 1.0, 0.66, and -0.83 respectively. The former two colors indicate a K-type main-sequence star. The negative (i-z) color is due to bad sky in the z-band images. Further, we find there are 439 sources fainter than or equal to the brightness of ASASSN-24fw. We define the chance alignment probability (P(X)) of a star within 3" radius of ASASN-24fw as:
where the FoV is 13.2 x 13.2 sq. arcmin and is the area of a circular region of 3" radius around ASASN-24fw. We find that P(X) is only 2% which indicates that there is a high chance that the companion might not be a chance alignment. But this can be only confirmed in future by conducting follow-up campaign to study the proper motion of this faint companion.
2.3 AAVSO
American Association of Variable Star Observers (AAVSO) (Kloppenborg, 2025) is a association of professional and amateur astronomers that observe variability related phenomena and catalog their high quality photometric observations. This association of professional and amateur astronomers regularly observes and analyses variable stars and their observers are located across the globe. We obtained the AAVSO observed data from their website111https://www.aavso.org/data-download/.
2.4 NEOWISE
The NEOWISE mission (WISE Team, 2020) utilized the Wide-field Infrared Survey Explorer spacecraft and a survey strategy similar to WISE mission (Wright et al., 2010) to scan the sky between the ecliptic poles in great semi-circles and nearly perpendicular to the Earth- Sun line. This survey strategy created a set of multiple independent exposures on each point of the sky over the 10 year period and had observed the region including ASASSN-24fw in the past decade. Thus, we could obtain and combine all the single exposure NEOWISE photometry which is an array of one observation every six months. We found no infrared variability of ASASSN-24fw in the NEOWISE lightcurve as all NEOWISE observations were taken outside the eclipse.
2.5 TESS
The Transiting Exoplanet Survey Satellite (TESS) is a satellite designed to search for transiting exoplanets among the brightest (and nearest) stars over most of the sky (Ricker et al., 2015). TESS orbits the Earth every 13.7 days on a highly elliptical orbit, scanning a sector of the sky spanning 24 x 96 deg2 for a total of two orbits, before moving on to the next sector. It captures images at cadences of 2 s (used for guiding), 20 s (for 1 000 bright astroseismology targets), 120 s (for 200 000) and 30 min (full frame images). The instrument consists of 4 CCDs each with a field of view of 24 x 24 deg2, with a wide band-pass filter from 600-1000 nm.
ASASSN-24fw has a TESS ID (TIC) = 262032775. It was detected in Sectors 7 and 33, with observations taken from 8 January 2019 to 1 February 2019 and from 17 December 2020 to 13 January 2021. This was the period when the star was bright and therefore we could obtain its lightcurve to search for periodicity. We did not find any peaks in the periodogram, suggesting that ASASSN-24fw is a slow rotator and has no significant intrinsic variability. Although ASASSN-24fw was again observed by TESS in December 2024, there is no lightcurve available due to its faintness.
2.6 Spectral Observations
We obtained a low resolution spectrum with Keck LRIS instrument (Oke et al., 1995; Rockosi et al., 2010) on 29 October 2024. The resolution of the grating element was 1300 and the exposure time was 2700s. We also obtained a second high resolution transmission spectrum using the HiRES instrument on 28 January 2025. The resolution of the grating element that was used was 25,000 and the exposure time was 600s. Both of these observations were conducted when the star was dimmed and were extracted using Pypelt (prochaska2020). The wavelengths were then calibrated using a combination of Th-Ar spectra and telluric absorption with the standard star observations using xtellcor (Vacca et al., 2003). We also obtained a NIR low resolution spectrum (R 3000) using the TripleSpec instrument (Herter et al., 2008) at the 5.1m Hale Telescope at Palomar Observatory on 21 December 2024. The spectrum was extracted using the IDL package spextool (Cushing et al., 2004). The extracted spectra were also flux calibrated and corrected for telluric absorption in a similar way.
2.7 Spectral Energy Distribution
We form the spectral energy distribution of ASASN-24fw using publicly available pre-dimming photometry from various surveys. For this, we obtained single epoch photometric data in the optical from SDSS (Almeida et al., 2023), PANSTARRS DR2 (Flewelling, 2018), Gaia DR3 (Gaia Collaboration et al., 2023), HST (Whitmore et al., 2016), and in the infrared 2MASS (Skrutskie et al., 2006) and WISE (Wright et al., 2010). We corrected these magnitudes (as well as our observed spectra during dimming) for extinction using the mean value of pre-dimming reddening from Schlafly & Finkbeiner (2011) reddening maps. These maps are updated from Schlegel et al. (1998) reddening maps and use the standard Fitzpatrick (1999) reddening law. We find a relatively low value of E(B-V) 0.05 for ASASSN-24fw and a clear presence of infrared excess in the WISE points as shown in the first panel of Figure (2). In the second panel of the Figure (2), we can also see the SED formed using the spectrum observed during the dimming. In the following sections, we describe in detail about our SED fitting methods.
2.8 Extinction calculations using LCOGT data
Our aim of obtaining multi-band LCOGT observations was to monitor the variation in color of ASASSN-24fw during dimming. The variation in color can then be used to compute the specific reddening vector and hence the composition of the occulting material. Unfortunately, our post-dimming baseline of ASASSN-24fw using LCOGT observations in the four SDSS filters is not very long. However, it is still enough to compute the reddening vector and hence constrain the primary composition of the occulting material. We generate color-color plots using g’, r’, i’, and z’s) bands and use optool (Dominik et al., 2021) to calculate the opacities (= +) values for a range of dust grain compositions and two size distributions. We used the built-in refractive indices from optool to generate the values for materials including astronomical silicate (Draine, 2003), various crystalline and amorphous iron-magnesium silicates (Dorschner et al., 1995; Jaeger et al., 1998; Fabian et al., 2001; Suto et al., 2006), carbon-bearing species (Draine & Malhotra, 1993; Henning & Stognienko, 1996), quartz (Kitamura et al., 2007), and iron particles (Henning & Stognienko, 1996).
We consider a log-normal grain distribution with a width of 0.5 dex and a free peak size, considering contributions from sizes between 0.1 and 1.1m in steps of 0.01m when calculating , determined from the complex refractive indices assuming Mie theory (Mie, 1908). For a Sun-like star, the typical blowout size for a dust grain is 0.5m (Krivov, 2010), so we include transient dust with this range of grain sizes. The optical depth and extinction of a dust cloud that fully covers the stellar disc are then calculated from the values for each combination of composition and size distribution, which are then compared to the measured reddening vectors by least-squares fit to identify those which best match the observations. For ASASSN-24fw, we find that no combination of compositions available in optool and grain size produces a reddening vector that fits the data. This indicates that the occulting material comprises of large bodies and may also contain gas.
3 Nature of ASASSN-24fw
| Parameter | Value |
| Gaia ID | Gaia DR3 3152916838954800512 |
| PANSTARRS ID | 115441063290606954 |
| ALLWISE ID | J070518.97+061219.4 |
| 2MASS ID | 07051897+0612195 |
| Parallax (mas) | 0.9583 0.0152 |
| (mas/yr) | -3.75 0.01 |
| (mas/yr) | -7.61 0.01 |
| (mas/yr) | 5.13 0.01 |
| (mas/yr) | -6.77 0.02 |
| Radial Velocity (km/s) | 36.87 3.60 km/s |
| (km/s) | 5.33 0.13 |
| (km/s) | -7.06 0.28 |
| (km/s) | 52.36 6.05 |
| (km/s) | -20.07 2.12 |
| (km/s) | 3.97 0.2 |
| ruwe | 0.97 |
| G | 12.823 0.003 |
| Ag | 0.16 0.03 |
| GBP | 13.066 0.008 |
| GRP | 12.430 0.006 |
| (GBP-GRP) | 0.636 0.001 |
| Teff | |
| log(g) | |
The Gaia DR3 (Gaia Collaboration et al., 2023) properties of ASASN-24fw are summarized in Table (1). Here we also show the PANSTARRS, 2MASS, and ALLWISE IDs. The G-band absolute magnitude was determined from the extinction () corrected apparent G magnitude by subtracting the distance modulus, which was calculated from the distance of 104417 pc derived from the trigonometric parallax. Here, we have also included the 3D peculiar velocity computed using the equations by Reid et al. (2009), and Poleski (2013). We find that ASASSN-24fw belongs to the Milky Way’s thick disc distribution. We use galpy (bovy2015), a Python package to compute the orbit of ASASSN-24fw in Milky Way in the past 50 Ma. galpy supports orbit integration in a variety of gravitational potentials, evaluating and sampling various distribution functions, and the calculation of action-angle coordinates for all static potentials. We find that the orbit of ASASN24fw does not intersect with any star formation region. Therefore this tells us that although ASASN-24fw has a persistent infrared excess, it is unlikely to be a Young Stellar Object (YSO) and is more likely to be on main-seqenuce.
3.1 Search for intrinsic variability
We searched for variability related to inherent stellar phenomena like star-spots, accretion, or rotation which is high in pre-main-sequnece (PMS) stars in the TESS lightcurve of ASASSN-24fw. We find a peak at 0.41 d with a very small false-alarm probability. But the phase-folded lightcurve about this period is flat indicating that the frequency may not have any real significance. Since we do not have TESS lightcurve during dimming episode, we could not check the during dimming periodogram. But, instead we used ASAS-SN lightcurve to find repeating signals in the lightcurve of both the pre- and during dimming phase. Here we saw a peak at 1 day in both the periodograms which is actually common in ground-based photometry, and does not reflect any variability related to the star. As YSOs have high magnetic field, accretion and star spot activity, our analysis re-iterates that ASASSN-24fw is unlikely to be on pre-main sequence and is in fact more evolved.
3.2 Fundamental stellar parameters using Stellar Evolutionary Models applied pre-dimming SED
In the pre-dimming SED, we clearly see an infrared excess particularly in the mid-infrared WISE data points (the first panel of Figure (2)). We fit this SED with Castelli-Kurucz models222We have used VOSA (VOSA v7.5, bayo2008) https://svo2.cab.inta-csic.es/theory/vosa/ to perform this analysis (Castelli & Kurucz, 2003) and find the best fitting model has [Teff,log(g),[Fe/H]] = [6250,5.0,-0.5]. This model is shown in cyan color in all the panels of Figure (2).
We then applied the Bayesian inference code of del Burgo & Allende Prieto (2016, 2018) on a grid of stellar evolution models constructed from the PARSEC v1.2S tracks (Bressan et al., 2012; Chen et al., 2014, 2015; Tang et al., 2014) to derive the fundamental stellar parameters of ASASSN-24fw. This choice is supported by the good statistical match obtained for detached eclipsing binaries, particularly for stars on the main-sequence, where the measured dynamical masses and corresponding predictions are consistent on average to within 4% (del Burgo & Allende Prieto, 2018).
The grid of PARSEC v1.2S models arranged in this code consisted of ages ranging from 2 to 13,800 Ma in steps of 5%, and [M/H] from -2.18 to 0.51 in steps of 0.02 dex. We adopted the photometric passband calibrations of Riello et al. (2021), and the zero points of the VEGAMAG system. We fed our code by the following three input parameters: the absolute magnitude , the color (also adjusted for reddening and extinction), and the iron to hydrogen abundance (assumed to be solar [Fe/H]= dex.
Due to the presence of inherent infrared excess, as expected for YSOs, we also adopted a prior in the code that ASASSN-24fw is on PMS and recomputed the fundamental stellar parameters. In this case, we find that all the parameters except age of the star remain similar. We find that if the star is on the PMS, its age is 7 Ma and mass is 1.470.05 M. Conversely, if it is on the main-sequence, its age is 2.75 Ga and mass is 1.690.08 M. The other parameters are similar to each other in the uncertainty range. These results are listed in Tables Table (2 and 3).
In Figure (2), we show the two different Kurucz models corresponding to the stellar atmospheric parameters that we just obtained. In addition to the cyan colored model that we discussed earlier, the green colored Kurucz model corresponds to (Teff, log(g), [Fe/H]) = (6000, 4.0, -0.5) and the red colored Kurucz model corresponds to (Teff, log(g), [Fe/H]) = (6500, 5.0, 0.0) i.e. the two models that we obtained by using the code of del Burgo & Allende Prieto (2016, 2018).
In the second panel, we show the SED of the star during dimming formed using our observed spectra. We see that all the three Kurucz models fit well to the continuum but the main difference comes while checking the fit to the peak of the SED especially Ca II HK absorption lines. In the same plot, we have also shown an M8 brown dwarf NIR spectrum template overplotted on top of the observed Palomar NIR spectrum. This NIR template spectrum is taken from the NIRSPEC Spectroscopic Brown Dwarf Survey (McLean et al., 2007) which has a collection of spectral templates for the M, L, and T type brown dwarfs. We can clearly see that the template matches the continuum and there is an excess infrared flux from the end of H-band. This hints towards the elevated levels of infrared excess during dimming perhaps caused by circum-planetary rings to the brown dwarf. Since we do not have any mid-infrared observations during dimming, we cannot compare the pre-dimming and post-dimming fluxes for mid-infrared wavelengths. Finally in the bottom-panel we show the residual flux after subtracting the three model spectra from the observed spectra.
3.3 Blackbody fit
The emission in the mid-infrared and longer wavelengths comes from the circum-stellar dust. This dust particles absorb the incident stellar flux and re-emit into the longer wavelength regions. Thus the nature of the excess flux can tell us about the nature of the circum-stellar material. We can assume that the dust particles act as perfect black bodies and model the deviant points corresponding to the excess flux in the SED using a solid blackbody model. In Figure (2), it is visible that the excess flux actually manifests from the 2MASS band. So instead of a single component, we fit a two-component blackbody model. The warm component of our best fitted models gives us an effective temperature of 1072K for the primary and 387K for the secondary colder component. An effective temperature of 1072K indicates a warm circum-stellar material within 0.2 au radius from the star which is well below the dust sublimating temperature while an effective temperature of 387K hints at a colder disc located beyond 1.0 au from the star may consist of larger and rocky bodies. The two component black-body model is shown as a black dashed line passing through the WISE points.
3.4 Spectral Analysis
The spectra that we have obtained were during the dimming episode of ASASN-24fw. Thus they are a mixture of the stellar spectra and the occulter absorption sprectra. It is therefore possible that, along with the continuum, the occulter may have affected the lines in the stellar spectra. For example, the variation in the Hydrogen series lines and the Ca II HK lines in Figure (3) in the LRIS spectrum taken on 31 October 2025 and HiRES spectrum taken on 28 January 2025 indicates that the gaseous component in the occulter is affecting the stellar spectra. Further, in the HiRES spectrum, we see that the age indicator Li line is not detected (or very weak at the level of noise), Fe, and Mg lines are moderately detected. These features as well as the shape and depth of the Ca II HK lines are normally expected for an F-type main-sequence star. Unlike YSOs, we do not see any emission line in our spectra. However, we see a "reversal" feature or shape in the H absorption line of the HiRES spectrum. Whether this is astrophysical or due to spectral noise is not very clear from the single spectrum. However, to find the strength of this line, we calculated the equivalent width of the absorption () and central emission () component of the H line. For this, we fitted a Lorentzian profile to the absorption line by masking the emission line corrected for barycentric motion. We find that 93Å and 7.8Å and the radial velocity of the emission line is -7.14 km/s.
| Parameter | Value |
| () | 0.958 0.015 |
| (mag) | 2.57 0.03 |
| (mag) | 0.73 0.04 |
| [Fe/H] | 0.00 0.20 |
| Effective temperature (K) | 6146 99 |
| Radius () | 2.28 0.08 |
| Mass () | 1.47 0.05 |
| Mean density () | 0.125 0.014 |
| Surface gravity (, cgs) | 3.89 0.03 |
| Luminosity (/) | 6.67 0.22 |
| Bolometric magnitude (mag) | 2.68 0.04 |
| Age (Ga) | 2.75 0.24 |
| Parameter | Value |
| Effective temperature (K) | 6076 121 |
| Radius () | 2.35 0.11 |
| Mass () | 1.69 0.08 |
| Mean density () | 0.132 0.013 |
| Surface gravity (, cgs) | 3.927 0.022 |
| Luminosity (/) | 6.74 0.23 |
| Bolometric magnitude (mag) | 2.67 0.04 |
| Age (Ma) | 7.0 1.8 |
4 Nature of The Occulter
4.1 Lightcurve Modelling
While we focused on the earlier on characterizing the star, now we focus on understanding the occulter. The shape of the lightcurve can tell us a lot about its nature eg. small and regular dips can be attributed to exoplanets, while the irregular episodes of dimming can be attributed to patchy or fragmented circum-stellar material. However, a long period and large scale dimming requires an object comparable to the size of the star. A single dark body comparable to the size of the star is physically impossible. But, the occulting body could be considered as a system of central body with massive circum-planetary rings and depending on the orientation of the ring system, a large portion of the star light can be blocked for longer duration. We proceed with lightcurve modelling in two steps; first we model the lightcurve using a parametric template that matches the shape of the lightcurve. This helps us to derive some basic parameters like the center of dimming, characteristic timescale, and lightcurve depth. In the next step, we can find a ring system geometry which can best explain the shape of the lightcurve. Here the information from the first model helps in setting the priors for Bayesian analysis used to maximize the likelihood of the fit. This is explained in the following sections.
4.1.1 Transit-Profile model
As the nearly symmetric lightcurve (see Figure (4)) looks like an inverse hat, we fit the transit profile of the lightcurve using a parametric Generalised Gaussian, defined by:
where () is the model brightness, is the maximum depth of the dimming, is the time of center of dimming, is the duration of the dimming, and is a parameter that defines the shape of the dimming. In case of ASASSN-24fw, should be high because of the flatness of the occulted lightcurve. ASASN-24fw dropped to 16.75 magnitudes during dimming which is close to the detection limit of 17.5 mag of ASAS-SN camera. Therefore, the lightcurve would be affected by noisy data points. In order to model the long-term trends in the lightcurve and to handle the noisy data points, in the next step we used the non-parametric Gaussian Process regression model. This was implemented through a Python package called george (Ambikasaran et al., 2015), with a squared exponential kernel (ExpSquaredKernel) given by:
where and are the kernel hyper-parameters. While the former hyper-parameter is the amplitude of the correlated noise in the whole lightcurve, the latter is the characteristic timescale of the noise in the whole lightcurve. Using these two hyper-parameters, george will try to find a mean model through the distributed data points. In this process, it can highlight any trends in the data which can otherwise be missed. Therefore, we combine these two models to capture the dimming and non-periodic trends in the lightcurve. We call our combined model as Transit-Profile model. We also performed the Bayesian analysis to maximize the likelihood of the fit and calculated the posterior distribution of our parameters. For this, we used the emcee (Foreman-Mackey et al., 2013) Python package which implements the Affine Invariant Markov Chain Monte Carlo (MCMC) Ensemble sampler algorithm introduced by Goodman & Weare (2010).
Our best-fitted Transit-Profile model is shown in the first panel of Figure (4) and the parameters with their 16th, 50th, and 84th percentile values listed in Table (4). This fit tells use few things; first, the star started dimming gradually 150 days (roughly in March 2024) much before the actual alert was issued during the ingress. The second phase of dimming started around 02 September 2024 continued up to 13.66 magnitudes around 16 September 2024. Finally, the third phase of ingress caused a drastic and sudden drop in magnitude and the brightness reduced to 16.70 around 02 October 2024. This indicates that there are multiple layers of varying densities and optical thickness of material around the occulting body, possibly hinting towards a gas-giant body with massive rings.
To check if our Transit-Profile model is really capturing various trends in the data or is just over-fitting the lightcurve. For this, we repeat the transit-profile model fitting on the ATLAS survey lightcurve. The depth of ATLAS camera is higher than the ASAS-SN camera and hence the data would be less noisy. We see that the wiggles or the "micro-trends" in the fit have reduced but the three phases or the "macro-trends" of ingress are similarly fitted (see the red and green colored data points and the model passing through them in the Figure (4)). This indicates that the long-term or the "macro-trends" in the lightcurve are really present and discovered using our method.
| Parameter | Value |
| 0.89 | |
| 137.73 | |
| (days) | 272.35 |
| (days) | 2460697.94 |
| 4.72 | |
| 683.94 |
| Parameter | Value |
| ()day-1 | |
| (HJD-2450000 days) |
4.1.2 Occulting disc model
We now understand that the dimming is not cause by a single solid body but by a system with large area, varying density and optical thickness. So in the next step, we try to model the lightcurve by a point like physical body having a massive circum-planetary disc similar to that of Aur type systems (for example, Rodriguez et al. (2015, 2016); Rappaport et al. (2019b); van der Kamp et al. (2022). We call this model as the Occulting disc model. To compute this model, we use the Python package named pyPplusS333https://github.com/EdanRein/pyPplusS (Rein & Ofir, 2023). In our model, the dimming is caused by elliptical projection of the massive circum-planetary rings on to the star. Unlike the short dips caused by transiting planets which cover a small portion of the stellar surface, the large scale discs block a significant portion of the stellar area and hence require a different approach to calculate the stellar brightness during dimming. pyPplusS’s approach is to use the Polygon-Segments algorithm which efficient and stable algorithm.
Typically, the parameters for computing a model in pyPplusS are: planet radius (Rp), the inner diameter of the ring (Rin), the outer diameter of the ring (Rout), the inclination angle of the ring from the orbital plane of the exoplanet (), the azimuthal angle by which the ring is rotated about the z-axis (), opacity (), and the limb-darkening coefficients of the star. In addition, the code requires time dependent coordinates of the exoplanet’s center and . We modified our modelling approach by considering a point like exoplanet surrounded by a massive ring system with sharp edges which is comparable to or larger than the stellar size. We further simplify our model by considering = (), and = dimensionless impact parameter ( 0). Here is the transverse velocity of the system which is assumed constant throughout the event, and is the time of the closest approach. Our simplification is made in a way that the exoplanet is traversing the star in an edge-on orbit and across its diameter. Thus our model parameters are: , , , , , and .
We explored this parameter space by looking for models that could produce the 275 day long transit event, while also minimizing the values. Here also we have made use of the Python package . Obviously, the models that would produce such long transits need to have low transverse velocities that likely indicate the large semi-major axis between the occulter and the primary star. Our best-fitted model is shown in the lower panel of Figure (4) and the parameters with their 16th, 50th, and 84th percentile values are shown in Table (5). We see that our model fits the lightcurve well, but does not accurately describe the lightcurve because of a couple of things. First, the assumption that we made that rings have sharp edges is not entirely correct, and second, the trajectory may not be traversing through the diameter of the star. Nevertheless, our model gives an estimate on the geometry of the system. We see that our planet is surrounded by massive rings with the outer diameter being roughly 16R or 0.17 au. This system is inclined to the orbital plane by nearly 40o, rotated about the rotational axis by nearly 50o and the transverse velocity of the occulter is about 1R per day or nearly 17 km/s.
As a check, we try another method to compute the transverse velocity of the occulter available in the literature. For example, van Werkhoven et al. (2014) calculate the transverse velocity based on the maximum gradient of the lightcurve. It requires the stellar linear limb-darkening parameter (), of the star. This parameter was computed using the jktld code of Southworth (2015). This code uses the stellar model parameters of the star to linearly interpolate the tables from Sing (2010) and calculates u for the required band passes. For ASASSN-24fw, we obtain u = 0.48 for the g’-band data of ASAS-SN survey. Using this value of , the steepest time gradient of the light curve, , and taking R= (from Table (2)), we get an estimate on :
For ASASSN-24fw, we get the maximum value of the gradient of the lightcurve is 4.970.43 10-6 Lday-1. Therefore, is 13.74 2.49 km/s. This is close to the orbital speed of 11.5 km/s derived from Kepler’s third law and assuming an orbital period of 16000 days (Nair & Denisenko, 2024) and a orbital semi-major axis of 17 au. This is also not very different from the velocity that we get from our lightcurve model. Now, we can estimate the radius of the occulter () = x /2 0.16 0.02 au. This is similar to the outer radius of the disc that we get from our model 0.17 au. Since for a stable disc, its radius should be smaller than the Hill radius of the secondary, the lower limit we get on the mass of the secondary is 3()3> 3.42 . Therefore, we can safely assume that the secondary is most compatible with a super-Jupiter or a brown dwarf with ultra-massive rings.
5 Discussion
5.1 Is ASASSN-24fw an R Corona Borealis (R CrB) type variable?
R CrB stars also show semi-regular pulsations with periods of multiple tens of days and amplitudes up to a few tenths of a magnitude (Percy, 2023); in contrast, as mentioned before, the star does not show significant variability in its TESS light curves. R CrB type variables come under the parent class of hydrogen-deficient stars, are found in a temperature range similar to that of ASASSN-24fw, and are known to show similar long (although sporadic) dimming behaviour (Clayton, 1996, 2012; Schaefer, 2024; Crawford et al., 2025). R CrB stars are also known to exhibit mid-IR excess due to warm circum-stellar dust shells (Chesneau et al., 2014; Rao & Lambert, 2015; Tisserand et al., 2020). The spectrum of R CrB stars are known for their hydrogen-less erratic and asymmetric dimming profiles and are dominated by mostly helium, nitrogen, and carbon enrichment in their atmospheres (see (Asplund et al., 2000; Pandey et al., 2021)). We found deep hydrogen absorption lines and no helium emission or P Cygni profile in our Keck optical spectrum (the second plot in Figure (2)). Moreover, no carbon absorption features or other spectroscopic R CrB signatures were found in our optical spectrum (see Pandey et al. 2021). As the available spectra were taken during the eclipse, we have not performed a detailed spectral analysis in search for other elements characteristic of hydrogen deficient stars (eg. Germanium Saini & Pandey 2025). However, this will be performed later when once we obtain a post-egress spectrum. Nevertheless it is highly unlikely that ASASN-24fw is a R CrB type of variable.
5.2 Is ASASSN-24fw a YSO?
YSOs are usually found near molecular clouds and often show optical variability and accretion signatures due to high magnetic fields and active circum-stellar zone comprised of an protoplanetary/proto-stellar material around them (Joy, 1945; Gahm et al., 2008; Zsidi et al., 2025). The circum-stellar material around a T Tauri type of variable star is divided in to two components: (1) the inner component, which is warmer and contains finer dust that accretes on to the host and (2) the outer component which is mixture of gaseous and solid material. Together they may manifest as two solid bodies causing infrared excess in the SED (eg. Varga et al. 2018 and references within). In the case of ASASSN-24fw, the persistent infrared excess is also fit by two black bodies.
Although the shape of the SED and the mass of ASASSN-24fw is similar to that of a T Tauri star, in the only pre-dimming spectrum that we have from Gaia DR3, no H emission line appears which may point towards any accretion activity. Similarly, in our low resolution spectrum obtained from Keck, we do not see any emission signatures. In our Keck high-resolution spectrum taken in January 2025, we see several deep hydrogen absorption lines. If the Ca II HK lines are unaffected by the occulter, then their strength resembles to that of an F-type main-sequence star. However, in the same spectrum, we see that the H absorption line is shallower unlike other Hydrogen series lines and has a shape. If the shape is not due to spectral noise then a component of the occulting material is Hydrogen rich and re-emitting the incident stellar radiation into H region. Moreover, the otherwise flat lightcurve of ASASSN-24fw only shows deep occultation every 43.8 years (Nair & Denisenko, 2024) that is likely caused by a large solid object. This is unlike any other lightcurve of a T Tauri star where we expect regular episodes of irregular dips.
In most cases the circum-stellar discs around YSOs are bright in X-ray and also radio sources owing to the strong magnetic fields and high temperatures from ongoing accretion activity (Feigelson & Montmerle, 1999; Espaillat et al., 2019). We searched through known X-ray catalogs like ROSAT 2RXS (Boller et al., 2016), XMMDR10 (Webb et al., 2022), XMMSL2 (XMM-SSC, 2018), Chandra CSC2 (Evans et al., 2024), and swift2SXPS (Kosiba et al., 2024) and found no pointing at the location of ASASSN-24fw. Thus we do not have evidence that the companion of ASASSN-24fw is a compact binary emitting X-rays. Given the lack of erratic photometric variability, no X-ray detections, and no evidence of narrow emission lines in the optical spectrum at the time of dimming, we do not believe that the dimming of ASASSN-24fw results from the activity in an accretion disc or that ASASSN-24fw is surrounded by a compact star. As mentioned earlier, the peculiar velocity of ASASSN-24fw indicates that it does not appear to have interacted with any star forming regions at least in past 50 Ma. Further, the peculiar velocity indicates that it belongs to the older population of stars in thick disc. Therefore the origin of the infrared excess is unlikely due to proto-planetary material.
5.3 Is the occultation caused by circum-stellar material?
As mentioned earlier, many young stars show the "dipping" behavior (Cody et al., 2014; Ansdell et al., 2016, 2019), i.e., multiple, repeated, short, and irregular dips in brightness. The most common known dipper stars are also young main sequence stars of F- or G- spectral type, but still accreting like classical T Tauri stars. For these accreting stars, Ansdell et al. (2016), and Nuñez et al. (2017) show a positive correlation between the depth of dips and the amount of infrared excess. In ASASSN-24fw we see an infrared excess as well as a deep but long dimming causing a flat base unlike the irregular shaped lightcurve by "dippers". The circum-stellar material causing infrared excess appears to have spread out rather than concentrated in clumps. This is supported by the fact that the mid-infrared NEOWISE lightcurve is also non-variable. Now the question arises whether this material accretes on to the star and is responsible for the ’w’ shaped H-alpha line that we see.
From the lightcurve we also see a phenomenon called as "forward-scattering", meaning the post-dimming color is not very different or slightly bluer than the original color of the star. This is possible when the material contains bodies that block the starlight but do not cause extinction effects. In terms of our Solar System, strong forward scattering has been observed in the Saturn B-ring system which is discussed by Hedman & Stark (2015). So the fact that we see forward scattering in the post-dimming lightcurve of the ASASSN-24fw could hint towards a dense mixture of dust and gas. This is also what we inferred from our dust composition analysis based on LCOGT data using optool. However, large icy bodies also cannot be ruled out. This composition with the large size of the disc suggests that the system is young ( 1 Ga), as the dust is usually depleted over time by radiation pressure, Poynting-Robertson drag, or accretion.
5.4 Nature and age of the system
The longer orbital period, depth of dimming and eclipse duration of ASASSN-24fw makes it a unique, and observationally challenging binary system. The crude stellar analysis based on our available data suggests that the star is unlikely to be a YSO although it has a persistent infrared excess. It is likely that the star has an active planetary environment and the catastrophic collisions between planetesimals has caused a debris disc around it. Although this can be verified, the distance of 1kpc would make it challenging to resolve the discs. Nevertheless, mid-infrared studies from JWST can tell us about the composition of the circum-stellar environment. Follow up series of observations from NIRCam or VLT SPHERE can infact be useful to measure the proper motion and companionship of the companion that we found in our LCOGT images.
The lower limit on the mass of the secondary that we derived indicates that the secondary is either a super-Jupiter or a brown dwarf. There is no clear distinction between the lower-limit of brown dwarf mass and upper limit of super-Jupiter objects, as they are mainly categorized based on the formation mechanisms. For example, the core-accretion theory (Pollack et al., 1996) can explain the formation of gas giant companions at separations 100 au, but gravitational instability or disc fragmentation (Boss, 1997) is required for heavier gas-giants. Although the number of planets formed by core accretion is more than those formed by gravitational instability, the increase in the detection of wide and massive planets as in case of ASASSN-24fw, suggests that both core-accretion and gravitational instability together could also be responsible for planet formation. This in fact requires further exploration and detecting and analyzing more such events would be a nice step in this direction.
One of the distinguishing factors between the super-Jupiters and brown dwarfs can be their discs. If the secondary is a super-Jupiter (with 5–13 MJup), it could host a large disc, provided it is young or resides in a debris-rich environment. Such a system could persist, especially at 20 au where dynamical clearing is slow (for example, Liu et al. (2023)). However, forming a super-Jupiter hosting circum-planetary disc with a size of 272 times the size of Jupiter is complicated as it can become gravitationally unstable owing to reduced lifetime. Instead, the companion being a more massive brown dwarf is more likely scenario.
Brown dwarfs are believed to form like stars via a gravitational collapse, and thus can retain the surrounding material (Stamatellos & Whitworth, 2009; Stamatellos & Herczeg, 2015), especially when the disc radius is well outside the Roche limit of the star. A disc of 0.20 au radius is consistent with observed discs around young brown dwarfs (e.g., in star-forming regions like Taurus Kurtovic et al. 2021). Such discs comprise of gas and fine dust (micron-sized grains) that scatter or absorb the stellar light effectively. This material is possibly replenished, from the ongoing accretion (Stamatellos & Herczeg, 2015; Zhu, 2015).
Although replenishing process may be responsible for the observed disc, this process is efficient only in the first few million years of the formation. As the system evolves, various "outside" forces come affect the process. Thus, if we consider that the star is an older main-sequence star, then an age of about 2 Ga introduces new challenges for the disc survival. On such longer timescales, the brown dwarf’s disc is expected to be depleted due to dynamical and radiative processes. Moreover, as a mature disc looses its opacity due to dust coagulation or dispersal, a significant luminosity drop in ASASN-24fw due to a older disc seems improbable. Although at larger separations, the brown dwarf’s disc would experience minimal stellar radiation, but the internal processes (e.g., viscous evolution) and external perturbations (e.g., passing stars) would likely deplete it long before 2 Ga. Therefore, a massive, opaque circum-planetary disc that is surviving for so long on its own is unlikely unless it is replenished by some process (e.g. by accretion from the circum-stellar environment which could be rich in material due to active collisions). Another process for the replenishment is the Bondi–Hoyle–Lyttleton accretion (Hoyle & Lyttleton, 1939; Bondi & Hoyle, 1944) which would have played a role if the brown dwarf would have passed through an unusually dense region of the ISM (e.g., a molecular cloud or diffuse gas) in the past.
5.5 Comparison with recent studies on ASASSN-24fw
After the event ended in mid-2025, it validated the predicted egress by Nair & Denisenko (2024). Soon thereafter, two groups Forés-Toribio et al. (2025) and Zakamska et al. (2025) reported their observations and conclusions. While both the studies agree on the large grain sized dusty component involved in the dimming event, and that ASASSN-24fw is an F-type star and evolved on a main-sequence, they differ on the interpretations of the event. But, Forés-Toribio et al. (2025) also find degeneracy in the age (and hence the mass of the star) like us. They also suggest the presence of a 0.25M M-dwarf secondary and that a circum-binary disc instead of a circum-planetary disc responsible for the dimming. Further, as they do not find any color change during eclipse (using their multi-band LCOGT and UKIRT observations), and find time-varying polarization during dimming, it was suggested that the occulting material is comprised of large (20m) carbonaceous or water ice grains rather than small silicate and CO2 grains.
Zakamska et al. (2025) undertook spectral study of the dimming and suggest that ASASSN-24fw is an evolved main sequence star and the dimming was caused by a truncated circum-planetary disc comprising of gas and dust of large sized grains. They provide mass of the occulter to be in between 0.5-19 MJ and the radius of the disc to be 0.7 au. This is similar to what we get from our lightcurve modeling. Moreover, their analysis using deep NaI D absorption lines which are in excess than that expected from stellar photosphere alone suggest the possibility of circum-stellar winds. Unlike their spectra neither we nor Forés-Toribio et al. (2025) found any metallic and H emission lines in our optical spectra. Interestingly, the optical spectrum of Forés-Toribio et al. (2025) taken on 9 October 2024 i.e. just after the star became dim, shows deep H absorption line, and our Keck LRIS spectrum taken on 29 October 2024 shows no H line. But our Keck HiRES spectrum taken on 25 January 2025 shows a H absorption line with a ’w’ shape or a central emission feature. Interestingly, Zakamska et al. (2025) find hydrogen emission lines in their Magellan and GHOST spectra taken on 3 February 2025 and 5 March 2025 respectively, after our Keck HiRES observations. As the only available pre-dimming spectra is from Gaia DR3 which also shows H absorption line, and that we have ruled out that ASASN-24fw is a YSO, indicates that the emission feature is unlikely to be coming from the star and maybe associated with the Hydrogen gas component in the occulting material which is re-emitting the absorbed stellar radiation into H region. The value of that we find in our HiRES spectrum is 8Å. This would have increased in the spectra by Zakamska et al. (2025) and indicates a transient nature of the H line.
5.6 Future plans
Long dimming events like ASASSN-24fw are open to interpretation just based on lightcurve and SED analysis. To improve our analysis, we plan to conduct post-dimming observations and determine the stellar spinning speed from the rotational broadening, metallicity and measure the Ca II H&K and Li elemental abundances to estimate the age and evolutionary status of the star.
We also plan to compare the out-of-eclipse spectra with the pre-dimming ones and get a transmission spectrum to further characterize the composition of the occulter. However, at this stage, it is evident that the nature of ASASSN-24fw is degenerated with its age and the occulter is massive in size and has rings made materials in varying densities. The out-of-eclipse spectra would also shed light on the whether the H line comes from the stellar photosphere and if ASASSN-24fw is a YSO with thick disc kinematics and away from any star-forming region.
We also plan to conduct out-of-dimming polarimetric study of ASASSN-24fw and nearby ISM to understand the amount of polarization and the nature of the circum-stellar environment causing the infrared excess. Similarly, the sub-millimeter observations will be helpful in detecting the geometry of the circum-stellar environment of ASASN-24fw. The JWST observations using MIRI will help us to detect whether the composition is dominated by the silicate or the carbonaceous dust. Infact if the survival of the massive circumplanetary disc is indeed due to Bondi–Hoyle–Lyttleton accretion, we could also see signatures of ISM molecules in the NIRSpec and MIRI spectrum. Finally, observations using high-angular resolution instruments on large telescopes would help us to check the companionship of the faint source detected in our LCOGT images. All the observations, would be helpful to create a complete picture about ASASSN-24fw and give us insight about the evolution of such systems, existence of gas-giants with massive rings and the environments in which they form. Systems like ASASSN-24fw would be of great importance for the upcoming extremely large telescopes.
6 Conclusions
We have analyzed the recent dimming event around the star ASASSN-24fw. It previously exhibited stable brightness over decades of observation and then abruptly started noticeable dimming on 02 September 2024. The star had a persistent mid-infrared excess pointing towards the presence of a circum-stellar material. The estimated fractional luminosity of this material is 0.5%. But this excess did not exhibit any periodic variation in the pre-dimming NEOWISE lightcurve suggesting that the material is spread out around the star and is not located in clumps. The blackbody fit requires two components with temperatures of 1072 K and 387 K. From our lightcurve modelling, we suggest that this material is unlikely to be responsible for the recent dimming event.
We modeled the lightcurve in two stages: (1) using a combination of Generalised Gaussian Distribution and Gaussian Process Regression called the Transit-Profile model and (2) assuming a point like planet with massive and inclined rings traversing the diameter of the star called the Occulting-disc model. Our first model suggests that the dimming actually started in March 2024 much before the actual alert was issued, and it progressed in stages. This indicates that the outer disc came in contact earlier than the main body much before the actual ingress. The different slopes of the model indicate that the density of the occulting material is varying and inversely proportional to the radial distance from the central body. While our model overfits the "micro-trends" due to the noise in the ASAS-SN lightcurve during the dimming, fitting a similar model to the ATLAS lightcurve suggests that the macro trends in the lightcurve representing various stages of dimming are significant. Our second model is a very simplified assumption of the occulter being a massive ring system around a gas-giant planet and passing across the stellar diameter. This model suggests that the inner diameter of the disc is 0.11 and the outer diameter of the disc is 16 or 0.17 au (within the hill sphere). Together this system weighs at least 3.42 and is having a constant transverse velocity of roughly 17 km/s. It also tells us that the ring system is has an opacity of 0.96 and is inclined to the orbital plane at an angle of 40o and azimuthally rotated about the z-axis at an angle of 50o. This orientation casts a massive elliptical shadow on the star that is responsible for the dimming event. This model does not fit the ingress very well because our assumptions that the planet is traversing the diameter of the star and the transverse velocity may actually be different for different layers or disc material. Further, our optool analysis using the post-dimming LCOGT data that we obtained from February to April 2025 indicates that the disc material is likely to be a mixture of gas and/or larger icy bodies.
We modeled the pre-dimming SED of the star using and also the Bayesian inference code developed by del Burgo & Allende Prieto (2016, 2018). Using the mean value of reddening towards the star, gives us an age of roughly 8 Ma, mass around 1.6 M and [Teff, log(g), [Fe/H]] = [6250, 5.0, -0.5]. In the latter code we consider two possibilities that the star is in the pre-main sequence phase, or on the main-sequence. Whilst most of the stellar parameters derived from these models are similar to what we get from VOSA, we get two different ages of 7 Ma and 2 Ga. Since this analysis is based on the pre-dimming photometry, an accurate age and composition of the star will also be determined from future high resolution spectroscopic observations. We will also compare this spectrum with the post-dimming spectrum that we already have to compare the elemental abundances and thereby gaining in-depth understanding on the composition of the occulter. While two models suggest that the star is young, we find that its kinematics belong to the thick disc distribution. It is not close to any star-forming region nor it has arrived from one in the past 50 Ma. We also check the TESS lightcurve of ASASN-24fw. If ASASSN-24fw is a YSO or a young star, then its high magnetic field or star spot activity should be reflected in the TESS lightcurve. But the lack of any variability indicates that ASASSN-24fw is a older main-sequence star.
We obtained two post-dimming spectra from Keck LRIS and Keck HiRES. While they resemble an F-type main-sequence star, we do not see any emission lines. On comparing the optical spectra of Forés-Toribio et al. (2025) and Zakamska et al. (2025) we find that the H emission feature started to appear after January 2025. So the conclusion on the evolutionary status of the star can be only confirmed using future out-of-eclipse high-resolution spectral observations.
Based on the massive ring size that we get, the companion is more likely to be a brown dwarf rather than a gas giant planet. As such massive circum-planetary discs should not survive beyond 100 Ma because the material would dissipate over time due to accretion, photo-evaporation, or dynamical clearing. Therefore, the existence of such a massive circum-planetary disc that is causing periodic dimming every 43.8 years means an extraordinary mechanism like a recent collisional replenishment, a anomalous stability, or Bondi–Hoyle–Lyttleton accretion processes are playing/might have playedrole. This challenges our understanding of circum-stellar disc lifetimes and brown dwarf environments and makes ASASN-24fw interesting system for future studies. If the age measured from future high-resolution spectra is consistent with the 2 Ga interpretation then the long-term existence of the circum-planetary scenario will be confirmed, and would require new theories about disc survival, debris disc dynamics, or unseen dynamical influences in mature systems.
Events like ASASSN-24fw are unique, open to interpretation and very important to understand stellar, circum-stellar disc, and planetary evolution. Future observations are therefore required to solve the puzzle of the evolutionary status of these systems.
Acknowledgements
We thank the referees for their suggestions and feedback to improve this manuscript. JPM acknowledges research support by the National Science and Technology Council of Taiwan under grant NSTC 112-2112-M-001-032-MY3. CdB acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant WEAVE: EXPLORING THE COSMIC ORIGINAL SYMPHONY, FROM STARS TO GALAXY CLUSTERS and the European Regional Development Fund (ERDF) with reference PID2023-153342NB-I00/10.13039/501100011033, as well as from a Beatriz Galindo Senior Fellowship (BG22/00166) from the MICIU. The Universidad de La Laguna (ULL) and the Consejería de Economía, Conocimiento y Empleo of the Gobierno de Canarias are also gratefully acknowledged for the support provided to CdB (2024/347). GH acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no. 695099). SZ acknowledges support from the Research Fellowship Program of the European Space Agency (ESA). SGD acknowledges generous support from the Ajax Foundation. This publication makes use of VOSA, developed under the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. VOSA has been partially updated by using funding from the European Union’s Horizon 2020 Research and Innovation Programme, under Grant Agreement No. 776403 (EXOPLANETS-A). This work has also made use of Astropy:444http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2022). We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research.
Data Availability
The spectroscopic data and LCOGT photometric data underlying this article will be shared on reasonable request to the corresponding author. The other photometric data used in this work are from public surveys.